OpenStudy (anonymous):

solve the equation for x: 2x - 5/7x = 36

OpenStudy (anonymous):

thank you so much! It sucks that my professor is making us do a prereq assignment when I signed up for the class to learn, not try to remember what I learned a few years ago ugh

OpenStudy (anonymous):

why did my responses disappear? i didnt get a chance to copy the equation :/

OpenStudy (anonymous):

 is the equation written like this

OpenStudy (anonymous):

no let me draw it real quick

OpenStudy (anonymous):

\[2x-\frac{ 5 }{ 7 }x=36\]

OpenStudy (anonymous):

2x-5/7x = 9/7x 9/7x = 36 x=28

OpenStudy (anonymous):

how did you get the 9?

OpenStudy (anonymous):

\[\frac{ 2 }{ 1 }x - \frac{ 5 }{ 7 }x--> \frac{ 14 }{ 7 }x -\frac{ 5 }{ 7 } x\]

OpenStudy (anonymous):

yeah i figured seconds before you replied lol. Thank you so much for taking the time to help!

Nnesha (nnesha):

 now solve for x